### v1

# Industrial Deposition of Wavelength–Shifting Films for Liquid Argon Photon Detection Systems

Abstract
The Deep Underground Neutrino Experiment (DUNE) Phase-II Far Detector is considering an unprecedentedly large-area ($\sim\!$2000 m2) photon detection system to achieve a target mean light yield of 180 PE/MeV. Meeting this requirement demands scalable, cost-effective, and high-quality wavelength-shifter (WLS) coatings capable of converting 127 nm liquid-argon scintillation light into visible photons with high efficiency. We report on the successful realization of an industrial physical vapor deposition (PVD) process for p-terphenyl (pTP) coatings, adapted from vacuum deposition techniques developed for OLED display manufacturing, to produce uniform WLS layers on large-area inorganic substrates, a task traditionally challenged by adhesion and uniformity issues at organic–inorganic interfaces. Surface characterization by profilometry and spectroscopic measurements confirms sub-10% thickness uniformity for 1-2 µm films and emission spectra consistent with high-quality pTP reference samples. The industrial process demonstrates reproducibility, scalability, and significantly reduced production time compared to laboratory-based methods, while maintaining the optical quality required for large-scale liquid-argon time-projection chamber (LArTPC) photon detection systems. These results establish a viable pathway for mass production of high-performance pTP coatings for DUNE FD3 and future neutrino experiments.
keywords:
Liquid argon detectors, Photon detectors, Wavelength shifters, Thin film deposition, Vacuum technology
1 Introduction
The Deep Underground Neutrino Experiment (DUNE) [2] is a next-generation long-baseline neutrino oscillation experiment designed to address fundamental questions in particle physics, including leptonic charge conjugation parity (CP) violation, the neutrino mass ordering, and proton decay. Its photon detection system (PDS) plays a critical role in detecting scintillation light from liquid-argon time-projection chambers (LArTPCs), enabling precise event timing, improved calorimetric energy reconstruction, and enhanced sensitivity to low-energy physics channels [18, 21].
In the DUNE Phase-I Far Detectors 1 and 2 (FD1/2) [2, 6], the baseline light-collection technology is the X-Arapuca light-trap design [8, 13]. This concept combines a primary wavelength-shifting (WLS) layer, an optical trapping mechanism with a secondary WLS layer, and silicon photomultiplier (SiPM) readout to achieve high photon-collection efficiency. Figure 1a illustrates the working principle of the X-Arapuca cell. In this configuration, the primary WLS is deposited directly onto a substrate facing the LAr volume, where it converts the 127 nm VUV scintillation photons into near-UV photons suitable for subsequent wavelength shift, light trapping, and detection.
For DUNE Phase-II Far Detector [1], the PDS design remains under active optimization, and several alternative light-trap geometries are being investigated. Preliminary optical simulations and prototype studies indicate that comparable photon-collection performance may be achievable even without a dichroic interface, motivating parallel development of simplified single-layer WLS coatings. The present work focuses on realizing such a coating through industrial-scale deposition of p-terphenyl (pTP) films on large-area substrates, evaluating their optical and mechanical performance as a potential solution for future DUNE Far Detector modules.
(a) X-Arapuca baseline geometry
(b) pTP absorption and emission spectra
Figure 1: (a) Working principle of an X-Arapuca cell, adapted from Ref. [8]. (b) Representative absorption and emission spectra of pTP. The blue curve shows the excitation light peaking at 266 nm, and the red curve shows the emission spectrum of pTP, reproduced from Ref. [3].
pTP is selected as the primary WLS material for its high quantum efficiency, fast re-emission time, and emission spectrum peaking in the near-UV ($\sim$340–360 nm), well matched to the subsequent optical elements [3] (see Fig. 1b). The quality of the pTP layer directly impacts the photon-detection efficiency through its conversion efficiency, emission-spectrum stability, and optical coupling to the trapping system.
Coating an organic wavelength-shifting material such as pTP onto inorganic substrates (e.g., borosilicate glass, quartz, or sapphire) presents intrinsic challenges. Differences in surface energy, thermal expansion coefficient, and chemical bonding properties often lead to poor adhesion, microcracking, or non-uniform film morphology. These effects are particularly pronounced when scaling to large-area substrates, where even slight variations in surface condition can result in substantial differences in film thickness and optical performance. Addressing these interface-related challenges is therefore essential for ensuring reproducibility and long-term stability of the coatings in large cryogenic detectors, and motivates detailed surface characterization of adhesion, morphology, and uniformity.
A variety of wavelength-shifting (WLS) coating techniques have been developed for large-area photodetector applications. Vacuum thermal evaporation provides precise control over film thickness and excellent optical quality, making it widely used for small- and medium-scale coatings; however, the throughput is traditionally limited for very large substrates [12, 4]. Spin coating can produce smooth and uniform layers on moderately sized substrates, although thickness uniformity and edge control tend to degrade as the coating area increases [5]. Doctor-blade coating offers a scalable approach suitable for large surfaces but involves direct mechanical contact, which may introduce surface defects or non-uniform regions if not precisely controlled [22]. Finally, solvent-based or spray-deposition techniques are simple and low-cost but typically require careful optimization of solvent choice, viscosity, and evaporation rate to achieve good adhesion and uniformity [16, 10]. Each of these methods has been used successfully in laboratory-scale experiments, but industrial-scale implementation demands a process that can deliver reproducibility, high optical quality, and scalability without sacrificing coating precision.
For the Phase-II Far Detector-scale deployment, the pTP coating must meet stringent requirements. First, it should have a thickness of order 1 µm (with an acceptable range of roughly 1–2 µm) to ensure efficient absorption of 127 nm light without unnecessary self-absorption or excess material, consistent with previous optimization studies of organic wavelength-shifting films [12, 4]. Second, the coating must achieve spatial uniformity at the $\lesssim 10\%$ level across each substrate, so that coating-induced light yield variations remain subdominant in the overall PDS response. Here, PDS response can be defined with minimum and average of the detector’s light yield [15]. Third, the films must exhibit robust adhesion and mechanical stability to withstand handling, cryogenic cycling, and long-term immersion in LAr. In addition, the chemical and optical purity of the pTP layer must be sufficiently high to preserve the characteristic emission spectrum and avoid quenching or spectral distortions from impurities or degradation products. Scaling production to the $\mathcal{O}$(2000 m2) required for the Phase-II Far Detector motivates a transition from manual laboratory setups to industrialized processes that deliver high throughput, reproducibility, and in-line quality control.
The following sections describe the industrial coating process and characterization methods (Sec. 2) and performance validation of pTP films on multiple substrate types (Sec. 3).
2 Methods
The coating procedure developed in this work was established through collaboration between Brookhaven National Laboratory (BNL) and LaserFiberOptics LLC, adapting industrial physical vapor deposition (PVD) processes commonly used in the optical and display industries for the production of high-uniformity organic thin films. The goal was to demonstrate that such an industrial approach can be directly applied to the deposition of pTP wavelength-shifting layers for large-area photon detection systems in liquid-argon detectors. Compared to traditional laboratory-based techniques, this process offers improved reproducibility, precise thickness control, and the throughput necessary for scaling to square-meter–level production. The following sections describe the deposition setup, characterization methods, and substrate studies carried out to establish the process parameters and verify coating performance. These methods together established a reproducible industrial coating workflow with full traceability from process parameters to optical performance.
2.1 Vacuum thermal deposition
PVD encompasses a class of thin-film techniques wherein material is deposited onto a substrate in a low-pressure environment. Reducing the ambient pressure to the millitorr range (or lower) increases the mean free path of vaporized species well beyond chamber dimensions, enabling largely ballistic transport from the source to the substrate and minimizing gas-phase scattering. This facilitates dense, uniform coatings and precise control of film thickness and composition.
In thermal evaporation, a source material is heated—via resistive heating, electron-beam bombardment, or related methods—until it vaporizes and subsequently condenses on cooler surfaces. Film properties depend on deposition parameters such as base pressure, substrate temperature, deposition rate, and source–substrate geometry. In the absence of a carrier gas, vacuum-deposited films often exhibit high purity and low defect density relative to solution-processed coatings. Widespread adoption in microelectronics and optics reflects advantages in scalability, batch-to-batch reproducibility, and compatibility with metals, dielectrics, and organic compounds.
2.2 Deposition procedure for pTP coatings
Pre-deposition substrate preparation
Following the conventional standard protocols of optical substrate preparation, the substrates are cleaned in ultrasonic baths using solvents followed by deionized water rinses. These steps were found to be essential for achieving high-yield, defect-minimized pTP films.
Plasma surface treatment
Plasma surface treatment was employed to enhance the adhesion and surface cleanliness of the substrates prior to pTP deposition. In this process, the substrate surfaces were exposed to a low-pressure reactive plasma that removes residual organic contaminants and modifies the surface chemistry, thereby improving bonding with the subsequently deposited pTP film. The plasma was generated from a mixture of N2, O2, and CF4 gases in a ratio of 2:10:1 and applied for approximately 15 minutes (Fig. 2). This gas mixture produces a chemically active plasma containing oxygen- and nitrogen-based species that promote surface activation, while CF4 contributes mild fluorination and surface cleaning effects commonly used in optical and semiconductor processing. The substrates were placed on metal meshes inside the plasma chamber to ensure uniform exposure.
Areas partially shielded by the supporting mesh exhibited reduced plasma exposure, resulting in lower pTP adhesion and thinner coating in those regions. To avoid such grid-pattern artifacts in the final coatings, all subsequent samples were positioned so that the active deposition surface was fully exposed to the plasma. This configuration provided consistent surface activation and significantly improved coating uniformity across the entire substrate area.
Figure 2: Plasma-treatment system and substrate placement. Mesh shadowing during treatment produced observable thickness modulation in early trials.
In situ ion bombardment and deposition parameters
The coater is equipped with an argon ion gun. Substrates mounted on a rotating dome were exposed to Ar ion flux for $\sim$10 minutes prior to deposition. This treatment improved film nucleation and mitigated isolated regions of poor wetting attributable to trace organics. The ion flux also modestly elevated the substrate temperature.
Deposition was performed at a working pressure of $1.6\times 10^{-4}$ Torr (turbo-molecular pumping). The pTP source boat temperature was maintained at $\sim$195 C, yielding a deposition rate of $\sim$5 Å/s. This temperature is well below the bulk boiling point ($\sim$389 C) and near the melting point ($\sim$212 C), as per standard pTP datasheets [11]. Operating in this regime was found to minimize spitting and produce uniform and optically smooth pTP films (see Fig. 3).
Figure 3: Uniform pTP film on a 143.75 mm $\times$ 143.75 mm B33 substrate.
pTP material and deposition campaigns
Three deposition campaigns were carried out as the process was optimized. The first campaign used a generic pTP powder on small witness samples of B33, quartz, and sapphire without plasma treatment, and was employed solely to explore basic PVD operating conditions. A second campaign used full-size 143.75 mm $\times$ 143.75 mm substrates (10 sapphire, 15 B33, and 10 quartz) coated to a nominal thickness of 1.5 µm. For each substrate type, approximately half of the samples received plasma surface treatment and the remainder were processed without it, still using generic pTP. This campaign informed the choice of plasma treatment, thickness range, and chamber conditions for the final recipe.
For the third campaign, the source material was upgraded to high-purity pTP from Thermo Fisher Scientific, and only B33 substrates were coated (15 samples at 1 µm thickness), all with plasma surface treatment applied. The change from generic to branded high-purity pTP led to noticeably cleaner emission spectra with a more clearly defined peak and higher apparent light yield. Unless otherwise stated, the B33 emission and thickness results presented in the analysis in Sec. 3 focus on these best-controlled samples: high-purity pTP on B33 from the third campaign, and plasma-treated quartz and sapphire from the second campaign.
Throughput and scaling for mass production
[…]
(a) Three-tier rotating dome
(b) Dome loaded prior to deposition
Figure 4: Throughput scaling with a rotating substrate dome.
In addition to the deposition step itself, the auxiliary process steps were evaluated for scalability. Substrate cleaning and handling scale linearly with batch size but are not rate-limiting relative to the PVD cycle. Plasma surface treatment is performed with substrates loaded in multiple layers within the plasma chamber, whose capacity exceeds that of the PVD coater; as a result, the nominal 15-minute plasma treatment time remains unchanged when operating at full coater load. Similarly, the in-situ argon ion bombardment step is applied simultaneously to all substrates mounted on the rotating dome and therefore retains a fixed duration of approximately 10 minutes independent of batch size. Consequently, the overall production throughput is dominated by the pTP deposition time, and the additional preparation steps do not introduce a scaling penalty for DUNE Phase-II FD-scale production.
2.3 Photo-spectrometer setup and measurement procedure
The emission properties of the coatings were characterized using a refurbished McPherson 234/302 VM monochromator/spectrometer equipped with a high-intensity deuterium lamp, providing excitation in the 110–550 nm range (Fig. 5). Initial measurements of fluorescence efficiency and UV transparency were performed at room temperature.
(a) Monochromator
(b) Schematic
Figure 5: Monochromator setup used for emission-spectrum measurements.
Emission spectra were recorded with a high-sensitivity cooled-CCD spectrometer (QE Pro 65000 by Ocean Optics). The deuterium lamp output is wavelength-selected by a motorized grating with an effective bandwidth of about 4 nm full width at half maximum (FWHM); the wavelength selection slit is fully opened to maximize excitation light intensity, and the light is then delivered to the sample chamber via a collimated beam path. The optical train operates in an argon-purged enclosure to minimize O2/H2O absorption at short wavelengths.
2.4 Profilometer measurements
To measure the thickness of the deposited pTP layer, a stylus profilometer was used. Because the instrument measures relative height differences, each substrate was prepared with a small uncoated region near one edge or corner to provide a distinct step between the bare substrate and the coated surface. This “step” was created by masking a narrow strip of the substrate during deposition so that pTP was not deposited in that area. The height difference between the coated and uncoated regions directly corresponds to the film thickness.
After deposition, multiple scans were taken across this masked step to quantify both the average coating thickness and its local variation. Scans typically covered a range of 3.5 mm from the uncoated edge into the coated region, with the central 0.5–2.5 mm interval used for uniformity analysis. The profilometer has a vertical resolution of approximately 2 nm, allowing precise measurement of small thickness variations across the surface. Together, the UV–Vis spectroscopic measurements and profilometer scans provided complementary surface characterization of the coated samples, enabling quantitative assessment of both optical and geometric uniformity.
2.5 Tested substrates
As summarized in Sec. 2.2, pTP coatings were deposited on borosilicate (B33) glass, quartz (fused silica), and sapphire substrates as part of the multi-campaign optimization program. Here we briefly summarize the relevant material properties; a comparison is given in Table 1.
B33 glass is a borosilicate glass widely used for optical and detector applications due to its low thermal expansion, chemical resistance, and relatively low cost. It is commercially available in large, uniform sheets and is straightforward to machine. Transmission in the near-UV is moderate, with increased absorption below $\sim$300 nm [19, 20]. Quartz (fused silica) provides excellent UV transmission to $\sim$160 nm and very low thermal expansion, supporting repeated cryogenic cycling [14, 9]. The primary limitation is cost for large-area panels. Sapphire (Al2O3) is extremely hard and durable, with broad UV–IR transparency and high thermal conductivity. It presents a stringent test for film adhesion and stability but is relatively expensive and challenging to process in large areas [17, 7]. Brief summaries and comparative properties are provided in Table 1.
Table 1: Comparison of tested substrates. UV cutoff and thermal-expansion values are approximate, compiled from manufacturer datasheets and standard references.
Substrate UV Transparency Cutoff Thermal Expansion Coefficient Mechanical Properties Relative Cost / Availability
[…]
3 Results
Where multiple deposition campaigns exist for the same substrate, we report the results from the best-controlled configuration as summarized in Sec. 2.2. The B33 emission spectrum and thickness measurement results shown here correspond to the optimized third batch, while the quartz and sapphire results are drawn from the plasma-treated subset of the second multi-substrate campaign.
3.1 Emission spectrum
Figure 6: Emission spectra measured from pTP-coated samples under UV excitation. The spectra are area normalized.
Emission spectra of pTP-coated samples were acquired with long integration times to resolve the near-UV emission band; a representative set is shown in Fig. 6. All emission spectra were measured using monochromatic excitation at 266 nm, selected by the monochromator from the deuterium lamp output. As a result, the spectral shape of the excitation source does not contribute to the measured emission spectra, and no unfolding of the lamp spectrum is required.
The measured emission profiles are in agreement with the canonical pTP spectrum (Fig. 1b), both in peak position (approximately 350 nm) and overall spectral shape. Spectra obtained from independent reference samples produced at Fermilab and Stony Brook University (SBU) exhibit the same features. The Fermilab reference sample consists of Thermo Fisher pTP deposited by PVD onto a B33 glass substrate, while the SBU reference sample uses Thermo Fisher pTP deposited by a solvent-based spray process onto an acrylic substrate. The close agreement among these spectra, despite differences in substrate and deposition technique, confirms that the industrially deposited films preserve the intrinsic optical behavior of pTP.
Peak variation among all measured samples was less than 5 nm, indicating strong reproducibility across substrates and deposition batches. This consistency demonstrates that neither substrate composition nor plasma pre-treatment measurably alters the photophysical response of pTP under the conditions studied here.
During earlier process-development runs using a generic pTP powder (see Fig. 7), we observed less well-defined emission features and lower apparent light yield, indicating that control of the pTP source purity is an important factor in achieving reliable, high-quality, and reproducible emission spectra. This observation is consistent with the purity requirements discussed in Sec. 1.
3.2 Thermal and cryogenic stress response
To assess the robustness of the wavelength-shifting films under realistic detector conditions, preliminary thermal and cryogenic stress tests were performed using B33 samples from the second deposition campaign (see Sec. 2.2). These samples, which employed a generic pTP source and included plasma-treated substrates, were used for initial robustness studies so that the final high-purity B33 batch could be reserved for detailed optical characterization.
Samples underwent multiple liquid-nitrogen cooling and warm-up cycles, after which no macroscopic delamination, peeling, or cracking was observed. Minor visual changes were noted on quartz and sapphire substrates during warm-up, possibly associated with surface condensation or transient moisture effects rather than degradation of the coating itself. Follow-up emission-spectrum measurements before and after thermal cycling are in progress to quantify any potential spectral shifts or changes in intensity. Initial observations, illustrated in Fig. 7, indicate that the emission profile remains stable within experimental uncertainty, both in spectral shape and light yield, suggesting that the industrial pTP coatings maintain optical integrity and are compatible with cryogenic operation in LArTPC photon detection systems.
Figure 7: Comparison of emission spectra before and after the cryogenic cycle.
3.3 Coating thickness and uniformity
Film thickness was measured by profilometry using a shallow step created by masking prior to deposition, so that the stylus scanned across a boundary between bare substrate and coated region. Multiple positions per sample were scanned to reduce bias from substrate non-flatness. As in Sec. 3.1, the B33 results shown here correspond to the optimized third deposition campaign using high-purity Thermo Fisher pTP, while the quartz and sapphire results are drawn from the plasma-treated subset of the second campaign (see Sec. 2.2).
(a) Thickness vs. lateral distance (B33)
(b) Thickness distribution (B33)
(c) Thickness vs. lateral distance (Sapphire)
(d) Thickness distribution (Sapphire)
(e) Thickness vs. lateral distance (Quartz)
(f) Thickness distribution (Quartz)
Figure 8: Profilometer results for representative B33, sapphire, and quartz substrates. Left: thickness vs. lateral scan distance, where the gray band indicates the lateral range used for thickness statistics; right: thickness distribution over 0.5–2.5 mm from the step edge.
Figure 8 illustrates typical scans. For sapphire, an apparent slope in the thickness vs. lateral distance is consistent with substrate bow rather than coating non-uniformity, complicating direct cross-sample comparisons. To standardize the metric, we report the standard deviation of thickness within the 0.5–2.5 mm interval. For each sample, six scans at different edge positions were taken; the replicate least affected by substrate flatness was selected for the summary in Table 2.
All coatings reached the target thickness range of 1–2 µm with less than 10% variation across the scanned regions. B33 glass consistently yielded the lowest standard deviations (40–60 nm), demonstrating excellent uniformity; quartz followed with <100<100 nm variation, and sapphire exhibited larger spread due to substrate bow and differing surface energetics. These results mark a significant improvement over early laboratory-scale coatings, which exhibited substantially larger thickness variations, indicating that the optimized industrial process achieves reliable large-area uniformity suitable for mass production. Such consistency across different substrates confirms the scalability and readiness of this approach for DUNE Phase-II Far Detector deployment from the coating-production standpoint.
Table 2: Representative film thickness statistics per sample and substrate. The replicate with minimal substrate-flatness bias (see text) is reported for each sample.
Sample number Substrate Mean thickness (nm) Std. dev. (nm)
S2 Sapphire 1460.6 118.1
S21 Quartz 2179.9 77.9
S29 Quartz 1372.9 99.4
S35 Sapphire 926.5 188.0
S45 B33 940.2 41.5
S46 B33 828.1 39.6
4 Conclusions
This work represents, to our knowledge, the first successful demonstration of industrial-scale physical vapor deposition of $p$-terphenyl coatings on multiple optical substrates, including borosilicate glass, quartz, and sapphire, with comprehensive surface characterization confirming that the films are smooth, uniform, and adherent, and meet the optical and mechanical quality requirements for photon detection systems in liquid argon time projection chambers such as DUNE Phase-II Far Detector. The industrial process yields consistent 1–2 µm coatings with sub-10% uniformity and has proven cost-effective and highly scalable; the demonstrated throughput indicates that the full 2000 m2 coating required for DUNE could be completed within approximately one year of production.
Beyond demonstrating large-scale manufacturability, this work addresses a long-standing materials challenge: achieving uniform, adherent organic coatings on inorganic substrates. The success of the industrial deposition process shows that interface control through optimized plasma and ion treatments can effectively overcome adhesion and uniformity issues that have traditionally limited such systems. This capability validates the robustness of the present approach and opens the path to broader applications of organic scintillators and wavelength-shifting films in cryogenic detector environments. Ongoing work focuses on further improving surface uniformity to enable next-generation light-trapper geometries with enhanced photon-collection efficiency.
Acknowledgements.
This work is supported by the Laboratory Directed Research and Development (LDRD) program at Brookhaven National Laboratory and by the U.S. Department of Energy, Office of Science, Office of High Energy Physics. The authors would also like to thank our industrial partner, LaserFiberOptics LLC for their dedicated collaboration and technical contributions.